3(5x-2)=54
How do i solve this? The solution is X=4 I don't understand how X can be 4.
-5 + X = -21 How do i solve this? The solution is -16. I don't understand how it can be -16
7-4X > 3-3X How do i solve this? The solution is < 4 I don't understand how the answer can be < 4
liobacasey16@yahoo.com I take on line class, have no text book.

Answer by checkley71(8403)

SUGGESTION: GO TO THE LIBRARY & GET A TEXT BOOK
------------------------------------------------
3(5X-2=54
15X-6=54
15X=54+6
15X=60
X=60/15
X=4 ANSWER.
PROOF
3(5*4-2)=54
3(20-2)=54
3(18)=54
54=54
------------------------------------------
-5+X=-21
X=-21+5
X=-16 ANSWER.
PROOF
-5-16=-21
-21=-21
-------------------------------------------
7-4X>3-3X
-4X+3X>3-7
-X>-4
X<4 (WHEN DIVIDING BY A NEGATIVE NUMBER (HERE IT IS -1) WE REVERSE THE > SIGN TO A < SIGN).
PROOF
LET X BE EQUAL TO 3
7-4*3>3-3*3
7-12>3-9
-5>-6 WHICH IS TRUE.